OpenStudy (anonymous):

So you know how Sin pi/6 = 1/2 right? So if a problem says Sin^2 pi/6 = 1/4 ?

OpenStudy (anonymous):

I mean, " does Sin^2 pi/6 = 1/4 ? "

OpenStudy (amistre64):

squaring both sides ... yeah

OpenStudy (psymon):

Yep, absolutely. Because sin^2(pi/6) is the same as: \[(\sin \frac{ \pi }{ 6 })^{2} \] Which then it becomes a sort of order of operations. Parenthesis first, then exponents.

OpenStudy (anonymous):

Cool, thanks :D

OpenStudy (psymon):

Yep yep.